OpenStudy (anonymous):

find the slope of line passing through the points (5,9) and (5,-2) A)-11 B)0 C)11 D) undefined

OpenStudy (anonymous):

d

OpenStudy (aaliyahfairgood):

m = y2-y1/x2-x1 m=(-2-5)/(5-5) m=-7/0 m=undefined your answer is D) undefined